OpenStudy (anonymous):

Find the distance between the points (-2, 4) and (4, -6). 2√2 2√(10) 2√(34)

OpenStudy (anonymous):

distance between (x1,y1) and (x2,y2) \[\sqrt{(x2-x1)^{2} + (y2-y1)^{2}}\]

OpenStudy (anonymous):

2 sqrt 10? @niksva

OpenStudy (anonymous):

2 sqrt 2

OpenStudy (anonymous):

it is 2 sqrt(34)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[\sqrt{(4- (-2))^{2}+ (-6-4)^{2}}\]

OpenStudy (anonymous):

\[\sqrt{(6)^{2} + (-10)^{2}}\] hope u can do now

OpenStudy (anonymous):

\[\sqrt{36+100}\] \[\sqrt{136}\] \[\sqrt{2*2*2*17}\] \[2\sqrt{34}\]